OpenStudy (anonymous):

x^2-9x-10=0. I got x=19 and I just want to get confirmed or be corrected please

OpenStudy (psymon):

Nope. And you should get 2 values for x as well. Do you have an idea how to factor this?

OpenStudy (anonymous):

apparently not

OpenStudy (jackellyn):

You need to find 2 numbers that add up to -9 and make the product -10. (x )(x )

OpenStudy (psymon):

You need 2 factors of -10 that add up to -9.

OpenStudy (anonymous):

-10 and 1

OpenStudy (psymon):

Right. Now just put those numbers with x as jackellyn did above.

OpenStudy (psymon):

(x + factor1)(x + factor2)

OpenStudy (jackellyn):

Make sure you keep your signs right

OpenStudy (anonymous):

(x-10) and (x+1)

OpenStudy (jackellyn):

and now you can solve each and you have your roots x-10=0 and x+1=0

OpenStudy (psymon):

Right ^_^

OpenStudy (anonymous):

10 and -1?

OpenStudy (jackellyn):

Yep

OpenStudy (anonymous):

so how do I right that as an answer?

OpenStudy (jackellyn):

x=10, x=-1 (10, -1) (x-10)(x+1) Depends on how your teacher wants it.

OpenStudy (anonymous):

Okay thanks a lot I have a few of those problems left that really helps :)